OpenStudy (anonymous):

Use an idenity to find the value of sin19degrees csc19degrees .do not use a calculator

OpenStudy (anonymous):

does it help to know that \(\csc(x)=\frac{1}{\sin(x)}\) ?

OpenStudy (anonymous):

yes so it would be 1/sin 19 right

OpenStudy (anonymous):

no , it would be \[\frac{\sin(19)}{\sin(19)}\]

OpenStudy (anonymous):

since you are multiplying by \(\sin(19)\)

OpenStudy (anonymous):

ok thanks which =1